OpenStudy (anonymous):

In the past 4 years, a sporting goods store had two yearly losses of $140,000 and $56,000 and two yearly profits of $110,000 and $60,000. What was the net profit or loss over 4 years? A. –$26,000 B. $26,000 C. $170,000 D. $196,000

OpenStudy (anonymous):

@butterflydreamer

OpenStudy (butterflydreamer):

losses = negative profits = positive So what do you think you need to do for this question?

OpenStudy (anonymous):

so -140,000 + 56,000 - 110,000 + 60,000

OpenStudy (anonymous):

right

OpenStudy (butterflydreamer):

not quite

OpenStudy (anonymous):

how then

OpenStudy (anonymous):

-140,000 + -56,000 + 110,000 + 60,000

OpenStudy (butterflydreamer):

look at your question: It says the losses are 140,000 and $56,000 So this means ( -140 000 ) and (- 56 000) and two yearly profits of $110,000 and $60,000. So this means ( + 110 000) and ( + 60 000) Now add them all together.... -140 000 - 56 000 + 110 000 + 60 000 = ?

OpenStudy (igreen):

Add up the losses and profits and subtract.

OpenStudy (igreen):

110000 + 60000 = ?

OpenStudy (igreen):

140,000 + 56000 = ?

OpenStudy (anonymous):

1.170000 2.196000

OpenStudy (butterflydreamer):

yep now subtract so... 170 000 - 196 000 = ?

OpenStudy (anonymous):

-26,000

OpenStudy (anonymous):

A

OpenStudy (igreen):

Yep, that's our answer.

OpenStudy (anonymous):

ok thanks guys

OpenStudy (butterflydreamer):

no problem ^_^